\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi
  
  \bibitem{Abdollahpour} M. R. Abdollahpour, R. Aghayaria, and M. Th. Rassias, Hyers-Ulam stability of associated Laguerre differential equations in a subclass of analytic functions, Journal of Mathematical Analysis and Applications, 437(2016), 605- 612.

\bibitem{AlsinaGer} C. Alsina and R. Ger, On some inequalities and stability results related to exponential function,
 	J. Inequal. Appl.,  2 (1998), 373-380.
 	
%\bibitem{BaiSim} D. Bainov, P. Simeonov, Integral Inequalities and
%Applications, Kluwer Academic Publishers, Boston, 1992

\bibitem{BrzCadCipl} Brzdek J., C\u{a}dariu L., Ciepli\'{n}ski K., Fixed Point Theory and the Ulam Stability, Hindawi Publishing Corporation, J. Funct. Spaces, 2014, Article ID 829419, http://dx.doi.org/10.1155/2014/829419

\bibitem{BrzdekPopaRasaXu}
 J. Brzdek, D. Popa, I. Rasa, B. Xu, Ulam Stability of Operators, Elsevier, 2018.

\bibitem{Brzdek} J. Brzdek, D. Popa, and Th. M. Rassias, Ulam Type Stability, Springer, 2019.


\bibitem{CimpeanPopa} D. S. Cimpean and D. Popa,
	 On the stability of the linear differential equation of higher order with constant coefficients,
	Appl. Math. Comput.,  217 (2010),  4141-4146.
	
 \bibitem{FilatovSha}
A.N. Filatov, L.V. Sharova, Integral inequalities and theory of nonlinear oscillations, Moscow, Nauka, 1976.

 \bibitem{Hyers} D. H. Hyers, On the stability of the linear functional equation,
	 Proc. Natl. Acad. Sci. USA, 27 (1941), 222-224.


\bibitem{Jung} S.-M. Jung, Hyers-Ulam-Rassias Stability of Functional Equations in Nonlinear Analysis, Springer, 2011.

\bibitem{Jung2} S. -M. Jung, C. Mortici, and M. Th. Rassias, On a functional equation of trigonometric type, Applied Mathematics and Computation, 252(2015), 294-303.

    \bibitem{Jung3} S. -M. Jung and M. Th. Rassias, A linear functional equation of third order associated to the Fibonacci numbers, Abstract and Applied Analysis, Volume 2014 (2014), Article ID 137468

\bibitem{Jung4} S. -M. Jung, D. Popa, and M. Th. Rassias, On the stability of the linear functional equation in a single variable on complete metric groups, Journal of Global Optimization, 59(2014), 165–171.

\bibitem{Jung1} S.-M. Jung,
	 Hyers-{Ulam} stability of linear differential equations of first order, {III},
	 J. Math. Anal. Appl., 311 (2005), 139-146.
	
\bibitem{Jung2}
S.-M. Jung, Hyers-{Ulam} stability of linear partial differential equations of first order,
 Appl. Math. Lett.  22 (2009) 70-74.

\bibitem{JungLee}
S.-M. Jung, K.-S. Lee, Hyers-{Ulam} stability of first order linear partial differential equations with constant coefficients,
 Math. Inequal. Appl. 10 (2) (2007) 261-266.

\bibitem{JungRassias1}
S.-M. Jung, Th.M. Rassias, Ulam's problem for approximate homomorphisms in connection with Bernoulli's differential equation,
Appl. Math. Comput.  187 (1) (2007) 223 - 227.

\bibitem{JungRassias2} S.-M. Jung and Th.M. Rassias, Generalized Hyers-Ulam stability of Riccati differential equation,
Math. Inequal. Appl. 11 (4)(2008) 777-782.

\bibitem{JungRassias3}
S.-M. Jung, B. Kim and Th.M. Rassias, On the Hyers-Ulam stability of a system of Euler differential equations of first order,
Tamsui Oxf. J. Math. Sci.   24 (4) (2008) 381-388.

\bibitem{Kannappan} Pl. Kannappan, Functional Equations and Inequalities with Applications, Springer, 2009,


\bibitem{Khan1} Khan H., Khan A., Abdeljawad T.,  Alkhazzan A.,  Existence results in Banach space for a nonlinear impulsive system, Adv. Differ. Equ., 2019, 18, https://doi.org/10.1186/s13662-019-1965-z

\bibitem{Khan2} Khan A.,  Gómez-Aguilar J.F., Khan T.S., Khan H.  Stability analysis and numerical solutions of fractional order HIV/AIDS model., Chaos, Solitons Fractals, 2019, 122, 119-128

\bibitem{Khan3} Khan H.,  Abdeljawad T.,  Aslam M., Khan R.A., Khan A.,  Existence of positive solution and Hyers-Ulam stability for a nonlinear singular-delay-fractional differential equation, Adv. Differ. Equ., 2019, 104, https://doi.org/10.1186/s13662-019-2054-z

\bibitem{Khan4} Khan H., Gómez-Aguilar J.F., Khan A., Khan T.S., Stability analysis for fractional order adecvation reaction diffusion system, Physica A: Statistical Mechanics and its Applications, 2019, 521, 737-751

\bibitem{Lee} Y. -H. Lee, S. -M. Jung, and M. Th. Rassias, Uniqueness theorems on functional inequalities concerning cubic-quadratic-additive equation, Journal of Mathematical Inequalities, 12(1)(2018), 43–61.

\bibitem{Lee2} Y. -H. Lee, S. -M. Jung, and M. Th. Rassias, On an n-dimensional mixed type additive and quadratic functional equation, Applied Mathematics and Computation, 228(2014), 13–16.

  \bibitem{LunguCiplea}
 N. Lungu, S.A. Ciplea, Ulam-{Hyers}-{Rassias} stability of pseudoparabolic partial differential equations,
Carpatian J. Math. 31 (2) (2015) 233-240.


\bibitem{LC}
 N. Lungu,  C. Craciun,  Hyers-Ulam-Rassias stability of a hyperbolic partial differential equations,
  International Scholarly Research Network Mathematical Analysis,  2012, Article ID 609754, 10 pages
http://dx.doi.org/10.5402/2012/609754.

\bibitem{LunguMar}
N. Lungu, D. Marian, Hyers-Ulam-Rassias stability of some quasilinear partial differential equations of first order,
 Carpatian J. Math. 35 (2) (2019) 165-170.


\bibitem{LunguPopa1}
N. Lungu, D. Popa, Hyers-{Ulam} stability of a first order partial differential equation,
 J. Math. Anal. Appl.  385 (2012) 86-91.

  \bibitem{LunguPopa2}
N. Lungu, D. Popa, On the Hyers-{Ulam} stability of a first order partial differential equation,
 Carpatian J. Math. 28 (2012) 77-82.

\bibitem{LunguPopa3}
N. Lungu, D. Popa, Hyers-{Ulam} stability of some partial differential equation,
 Carpatian J. Math. 30 (2014) 327-334.

  %\bibitem{LunguRus}
%N. Lungu, I.A. Rus, On a functional Volterra-Fredholm integral equation, via Picard operators,
 %J. Math. Inequal. 3 (4) (2009) 510-527.

 \bibitem{MarianCipleaLungu}
 D. Marian, S.A. Ciplea, N. Lungu, Ulam-Hyers stability of a parabolic partial differential equation, Demonstr. Math. (to appear)

\bibitem{Milovanovic} G. V. Milovanovi\'c and M. Th. Rassias (eds.), Analytic Number Theory, Approximation Theory and Special Functions, Springer, 2014.
    
    \bibitem{Mortici} C. Mortici, S. -M. Jung, and M. Th. Rassias, On the stability of a functional equation associated with the Fibonacci numbers, Abstract and Applied Analysis, Volume 2014 (2014), Article ID 546046.

  \bibitem{NgocThLo}
L.T.P. Ngoc, T.M. Thuyet, N.T. Long,  On nonlinear Volterra-Hammerstein integral equation in three variables,
Nonlinear Functional Analysis and Applications  19 (2) (2014) 193-211.

 \bibitem{Obloza} M. Obloza,  Hyers stability of the linear differential equation,
	 Rocznik Nauk-Dydakt. Prace Mat.,  13 (1993), 259-270.

\bibitem{Park} C. Park and M. Th. Rassias, Additive functional equations and partial multipliers in  C*-algebras, Revista de la Real Academia de Ciencias Exactas, Serie A. Matemáticas, 113(3)(2019), 2261-2275.

 \bibitem{PopaPugna} D. Popa and G. Pugna,
	 Hyers-{Ulam} stability of {Euler}'s differential equation,
 	 Results Math., 69 (2016), 317-325.

\bibitem{PopaRasa1} D. Popa and I. Rasa,
	 On the {Hyers}-{Ulam} stability of the linear differential equation,
	 J. Math. Anal. Appl., 381 (2011), 530-537.

\bibitem{PopaRasa2} D. Popa and I. Rasa,
	 Hyers-{Ulam} stability of the linear differential operator with non-constant coefficients,
	 Appl. Math. Comput., 219 (2012), 1562-1568.

  \bibitem{PraRassias}
  A. Prastaro, Th. M. Rassias,  Ulam stability in geometry of PDE's,
 Nonlinear Functional Analysis and Applications  8 (2)(2003) 259 - 278.

  \bibitem{RezRassias}
 H. Rezaei, S.-M. Jung and Th. M. Rassias, Laplace transform and Hyers-Ulam stability of linear differential equations,
 J. Math. Anal. Appl., 403 (1) (2013), 244-251.

\bibitem{RL} I.A. Rus,  N. Lungu,  Ulam stability of a nonlinear hyperbolic partial differential equations, Carpatian J. Math. 24 (2008) 403-408.

\bibitem{Rus2} I.A. Rus,  Remarks on Ulam stability of the operatorial equations, Fixed Point Theory 10 (2009) 305-320.

\bibitem{Takahasi} S. E. Takahasi, H. Takagi, T. Miura and S. Miyajima,
	 The {Hyers}-{Ulam} stability constant of first order linear differential operators,
	 J. Math. Anal. Appl.,  296 (2004), 403-409.
	
\bibitem{Ulam} S. M. Ulam, A Collection of Mathematical Problems. Interscience, New York, 1960.
\end{thebibliography}

\documentclass[review,3p,times,12pt]{elsarticle}

\usepackage{amssymb}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{pifont}
\usepackage[colorlinks]{hyperref}
\usepackage{amsthm}
\newtheorem{theorem}{Theorem}[section]

\newdefinition{defn}[thm]{Definition}
\newdefinition{rmk}[thm]{Remark}
\newdefinition{ex}[thm]{Example}
\newproof{pf}{Proof}

\journal{Approximation and Computation in Science and Engineering , edited by N.J. Daras and Th.M. Rassias, Springer}

\begin{document}

\begin{frontmatter}

\title{On Hyers-Ulam-Rassias  stability of a Volterra-Hammerstein functional  integral equation}



\author{Sorina Anamaria Ciplea, Nicolaie Lungu, Daniela Marian, Themistocles M. Rassias}

\address{Sorina Anamaria Ciplea, Technical University of Cluj-Napoca, Department of Management and Technology, 28 Memorandumului Street, 400114, Cluj-Napoca, Romania\\
 {sorina.ciplea@ccm.utcluj.ro}}

\address{Nicolaie Lungu, Technical University of Cluj-Napoca, Department of Mathematics, 28 Memorandumului Street, 400114, Cluj-Napoca, Romania\\
{nlungu@math.utcluj.ro}}

\address{Daniela Marian, Technical University of Cluj-Napoca, Department of Mathematics, 28 Memorandumului Street, 400114, Cluj-Napoca, Romania\\ {daniela.marian@math.utcluj.ro}}

\address{Themistocles M. Rassias, National Technical University of Athens, Department of Mathematics, Zografou Campus, 15780 Athens, Greece\\ {trassias@math.ntua.gr}}

\begin{abstract}
The aim of  this paper is to study Hyers-Ulam-Rassias stability for a Volterra-Hammerstein functional  integral equation in three variables via Picard operators.
\end{abstract}

\begin{keyword}
 Volterra-Hammerstein functional  integral equation; Hyers-Ulam-Rassias  stability

\noindent {\bf 2010 Mathematics Subject Classification}: 26D10; 34A40; 39B82; 35B20

\end{keyword}

\end{frontmatter}

\section{1. Introduction}
Ulam stability is an important concept in the theory of functional equations. The origin of Ulam stability theory was an open problem formulated by Ulam, in 1940, concerning the stability of homomorphism \cite{Ulam}.
The first partial answer to Ulam's question came within a year, when Hyers \cite{Hyers} proved a stability result, for additive Cauchy equation in Banach spaces.
The first result on  Hyers-Ulam stability of differential equations was given by Obloza  \cite{Obloza}. Alsina and Ger investigated the stability of  differential equations $y'=y$ \cite{AlsinaGer}.
The result of Alsina and Ger were extended by many authors  \cite{CimpeanPopa}, \cite{Jung1}, \cite{JungRassias1}, \cite{JungRassias2}, \cite{JungRassias3}, \cite{PopaPugna}, \cite{PopaRasa1}, \cite{PopaRasa2}, \cite{RezRassias}, \cite{Rus2}, \cite{Takahasi}  to the stability of the first order linear differential equation and linear differential equations of higher order. For a broader study of Hyers-Ulam stability for functional equations the reader is also referred to the following books and papers: \cite{Abdollahpour}, \cite{Brzdek}, \cite{Jung}, \cite{Jung2}, \cite{Jung3}, \cite{Jung4}, \cite{Kannappan},  \cite{Lee}, \cite{Lee2}, \cite{Milovanovic}, \cite{Mortici}, \cite{Park}, \cite{Ulam}.

The first result proved on the Hyers-Ulam stability of partial differential equations
is due to A. Prastaro and Th.M. Rassias \cite{PraRassias}. Some results regarding Ulam-Hyers stability of partial differential equations were given by S.-M. Jung \cite{Jung2},
S.-M. Jung and K.-S. Lee \cite{JungLee}, N. Lungu and S.A Ciplea \cite{LunguCiplea}, N. Lungu and D. Popa \cite{LunguPopa1}, \cite{LunguPopa2}, \cite{LunguPopa3}, N. Lungu,  C. Craciun \cite{LC}, N. Lungu and D. Marian \cite{LunguMar}, D Marian, S.A. Ciplea and N. Lungu \cite{MarianCipleaLungu}, I.A. Rus and N. Lungu \cite{RL}.
 In \cite{BrzdekPopaRasaXu} Brzdek, Popa, Rasa and Xu presented a unified and systematic approach to the field.
 Some recent results regarding stability analysis and their applications were established by H. Khan, A. Khan, T. Abdeljawad and  A. Alkhazzan \cite{Khan1}, A. Khan,  J.F. Gómez-Aguilar, T.S. Khan and H. Khan \cite{Khan2}, H. Khan,  T. Abdeljawad,  M. Aslam, R.A. Khan and A. Khan  \cite{Khan3}, H. Khan, J.F. Gómez-Aguilar, A. Khan and T.S. Khan \cite{Khan4}. Results regarding fixed point theory and the Ulam stability can be found in \cite{BrzCadCipl}.

In this paper we consider the following Volterra-Hammerstein functional
integral equation in three variables%
\begin{align}\label{eq1}
&u\left( x,y,z\right) =g\left( x,y,z,h\left( u\right) \left( x,y,z\right)\right) \\
&+\int_{0}^{x}\!\!\! \int_{0}^{y}\!\!\! \int_{0}^{z}K\left( x,y,z,r,s,t,f_{1}\left(
u\right) \left( r,s,t\right) \right) drdsdt+ \int_{0}^{\infty
}\!\!\! \int_{0}^{\infty }\!\!\! \int_{0}^{\infty }F\left( x,y,z,r,s,t,f_{2}\left(
u\right) \left( r,s,t\right) \right) drdsdt \nonumber
\end{align}
via Picard operators.

The present paper is motivated by a recent paper \cite{NgocThLo} of L.T.P. Ngoc, T.M.
Thuyet and N.T. Long in which is studied the existence of asymptotically
stable solution for a Volterra-Hammerstein integral equation in three
variables.
Equation \eqref{eq1} is a generalization of equation (1.1) from \cite{NgocThLo}.

\section{2. Existence and uniqueness}

In what follows we consider some conditions relative to equation \eqref{eq1}.

Let $\left( E,\left\vert \cdot \right\vert \right) $ be a Banach space and $%
\Delta =\left\{ \left( x,y,z,r,s,t\right) \in \mathbb{R}_{+}^{6}:s\leq x,r\leq
y,t\leq z\right\} $. Let $\tau >0$ and the set%
\begin{equation*}
X_{\tau }:=\left\{ u\in C\left( \mathbb{R}_{+}^{3},E\right) \mid \exists M\left(
u\right) >0:\left\vert u\left( x,y,z\right)\right\vert  e^{-\tau \left( x+y+z\right)
}\leq M\left( u\right), \forall \left( x,y,z\right) \in \mathbb{R}_{+} \right\} .
\end{equation*}%
On $X_{\tau }$ we consider Bielecki's norm%
\begin{equation*}
\left\Vert u\right\Vert _{\tau }:=\underset{x,y,z\in R_{+}}{\sup }\left(
\left\vert u\left( x,y,z\right)\right\vert  e^{-\tau \left( x+y+z\right) }\right) .
\end{equation*}%
It is clear that $\left( X_{\tau },\left\Vert \cdot \right\Vert _{\tau
}\right) $ is a Banch space. In what follows we assume, relative to \eqref{eq1},
the conditions:
\begin{itemize}
\item[(C1)]
 $g\in C\left( \mathbb{R}_{+}^{3}\times E,E\right) ,K\in C\left(
\Delta \times E,E\right) ,F\in C\left( \Delta\times E,E\right) ,h\in
C\left( X_{\tau },X_{\tau }\right) ,f_{1}\in C\left( X_{\tau },X_{\tau
}\right) ,f_{2}\in C\left( X_{\tau },X_{\tau }\right) ;$

\item[(C2)]  there exists $l_{h}>0$ such that%
\begin{equation*}
\left\vert h\left( u\right) \left( x,y,z\right) -h\left( v\right) \left(
x,y,z\right) \right\vert \leq l_{h}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left(
x+y+z\right) },\forall x,y,z\in \mathbb{R}_{+},\forall u,v\in X_{\tau };
\end{equation*}

\item[(C3)] there exists $l_{g}>0$ such that%
\begin{equation*}
\left\vert g\left( x,y,z,e_{1}\right) -g\left( x,y,z,e_{2}\right)
\right\vert \leq l_{g}\left\vert e_{1}-e_{2}\right\vert ,\forall x,y,z\in
\mathbb{R}_{+},\forall e_{1},e_{2}\in E;
\end{equation*}

\item[(C4)] there exists $l_{K}\in C\left( \Delta , \mathbb{R}_{+}\right) $
such that%
\begin{equation*}
\left\vert K\left( x,y,z,r,s,t,e_{3}\right) -K\left(
x,y,z,r,s,t,e_{4}\right) \right\vert \leq l_{K}\left( x,y,z,r,s,t\right)
\left\vert e_{3}-e_{4}\right\vert ,
\end{equation*}
$\forall \left( x,y,z,r,s,t\right) \in
\Delta ,\forall e_{3},e_{4}\in E;$
\item[(C5)] there exists $l_{F}\in C\left( \Delta, \mathbb{R}_{+}\right) $
such that%
\begin{equation*}
\left\vert F\left( x,y,z,r,s,t,e_{5}\right) -F\left(
x,y,z,r,s,t,e_{6}\right) \right\vert \leq l_{F}\left( x,y,z,r,s,t\right)
\left\vert e_{5}-e_{6}\right\vert ,
\end{equation*}
$\forall \left( x,y,z,r,s,t\right) \in
\Delta ,\forall e_{5},e_{6}\in E;$
\item[(C6)] there exists $l_{f_{1}}>0$ and $l_{f_{2}}>0$ such that%
\begin{equation*}
\left\vert f_{1}\left( u\right) \left( r,s,t\right) -f_{1}\left( v\right)
\left( r,s,t\right) \right\vert \leq l_{f_{1}}\left\vert u\left(
r,s,t\right) -v\left( r,s,t\right) \right\vert ,\forall \left( r,s,t\right)
\in \mathbb{R}_{+}^{3},\forall u,v\in X_{\tau },
\end{equation*}%
\begin{equation*}
\left\vert f_{2}\left( u\right) \left( r,s,t\right) -f_{2}\left( v\right)
\left( r,s,t\right) \right\vert \leq l_{f_{2}}\left\vert u\left(
r,s,t\right) -v\left( r,s,t\right) \right\vert ,\forall \left( r,s,t\right)
\in \mathbb{R}_{+}^{3},\forall u,v\in X_{\tau };
\end{equation*}

\item[(C7)] there exists $l_{1}>0$ and $l_{2}>0$ such that%
\begin{equation*}
\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{f_{1}}l_{K}\left( x,y,z,r,s,t\right)
e^{\tau \left( r+s+t\right) }drdsdt\leq l_{1}e^{\tau \left( x+y+z\right)
},\forall \left( x,y,z,r,s,t\right) \in \Delta ,
\end{equation*}%
\begin{equation*}
\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }l_{f_{2}}l_{F}\left(
x,y,z,r,s,t\right) e^{\tau \left( r+s+t\right) }drdsdt\leq l_{2}e^{\tau
\left( x+y+z\right) },\forall \left( x,y,z,r,s,t\right) \in \Delta ;
\end{equation*}

\item[(C8)] $l_{g}l_{h}+l_{1}+l_{2}<1$;

\item[(C9)] \begin{align*}
& \left\vert g\left( x,y,z,h\left( u\right) \left(
x,y,z\right) \right) \right\vert
+\int_{0}^{x}\!\!\int_{0}^{y}\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt\\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert
F\left( x,y,z,r,s,t,f_{2}\left( 0\right) \left( r,s,t\right) \right)
\right\vert drdsdt\leq \alpha \exp \left( \tau \left( x+y+z\right) \right) ,
\end{align*}
$\forall \left( x,y,z,r,s,t\right) \in \Delta $;
\item[(C10)] there exists $m>0$ such that
\begin{equation*}
\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\left[ l_{f_{1}}l_{K}\left(
x,y,z,r,s,t\right) +l_{f_{2}}l_{F}\left( x,y,z,r,s,t\right) \right]
drdsdt\leq m,\forall \left( x,y,z,r,s,t\right) \in \Delta .
\end{equation*}
\end{itemize}

\begin{theorem}
Under the conditions $\left( C1\right) -\left( C9\right) $ the equation \eqref{eq1}
has in $X_{\tau }$ a unique solution $u^{\ast }.$
\end{theorem}

\begin{proof}
We consider the operator
\begin{equation*}
A:X_{\tau }\rightarrow X_{\tau },A\left( u\right) \left( x,y,z\right) :=%
\text{ second part of \eqref{eq1}}.
\end{equation*}%
First we prove that $A(u)$ maps $X_{\tau }$ in $X_{\tau }.$ For $u\in X_{\tau }$
we have:%
\begin{align*}
&\left\vert A\left( u\right) \left( x,y,z\right) \right\vert  \leq
\left\vert g\left( x,y,z,h\left( u\right) \left( x,y,z\right) \right)
\right\vert +\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\! \int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right) \right\vert
drdsdt \leq \left\vert g\left( x,y,z,h\left( u\right) \left( x,y,z\right)
\right) \right\vert  \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right) -K\left(
x,y,z,r,s,t,f_{1}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right) -F\left(
x,y,z,r,s,t,f_{2}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt .
\end{align*}
We obtain
\begin{align*}
&\left\vert A\left( u\right) \left( x,y,z\right) \right\vert
\leq \left\vert g\left( x,y,z,h\left( u\right) \left( x,y,z\right) \right)
\right\vert +\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( 0\right) \left( r,s,t\right) \right) \right\vert
drdsdt \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
\left\vert f_{1}\left( u\right) \left( r,s,t\right) -f_{1}\left( 0\right) \left( r,s,t\right) \right\vert drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }l_{F}\left(
x,y,z,r,s,t\right) \left\vert f_{2}\left( u\right) \left( r,s,t\right) -f_{2}\left( 0\right) \left( r,s,t\right)
\right\vert drdsdt \\
&\leq \alpha \exp \left( \tau \left( x+y+z\right) \right)
+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
l_{f_{1}}\left\vert u\left( r,s,t\right) \right\vert e^{\tau \left(
r+s+t\right) }\cdot e^{-\tau \left( r+s+t\right) }drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\left\vert u\left( r,s,t\right) \right\vert
e^{\tau \left( r+s+t\right) }\cdot e^{-\tau \left( r+s+t\right) }drdsdt \\
&\leq \alpha \exp \left( \tau \left( x+y+z\right) \right) +\left\Vert
u\right\Vert _{\tau }\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left(
x,y,z,r,s,t\right) l_{f_{1}}e^{\tau \left( r+s+t\right) }drdsdt \\
&+\left\Vert u\right\Vert _{\tau }\int_{0}^{\infty }\!\!\!\int_{0}^{\infty
}\!\!\!\int_{0}^{\infty }l_{F}\left( x,y,z,r,s,t\right)l_{f_{2}} e^{\tau \left(
r+s+t\right) }drdsdt
\end{align*}%
We have%
\begin{equation*}
\left\vert A\left( u\right) \left( x,y,z\right) \right\vert \leq \left[
\alpha +\left\Vert u\right\Vert _{\tau }\left( l_{1}+l_{2}\right) \exp \left( \tau \left( x+y+z\right)
\right) \right] ,
\end{equation*}%
hence $A(u)\in X_{\tau }.$

The operator $A$ is a contraction in $X_{\tau }$ with respect to $\left\Vert
\cdot \right\Vert _{\tau }.$ Indeed, for $u,v\in X_{\tau },$  we have:%
\begin{align*}
&\left\vert A\left( u\right) \left( x,y,z\right) -A\left( v\right) \left(
x,y,z\right) \right\vert  \leq \left\vert g\left( x,y,z,h\left( u\right)
\left( x,y,z\right) \right) -g\left( x,y,z,h\left( v\right) \left(
x,y,z\right) \right) \right\vert   \notag \\
&+\int_{0}^{x\!\!\!}\int_{0}^{y}\!\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right) -K\left(
x,y,z,r,s,t,f_{1}\left( v\right) \left( r,s,t\right) \right) \right\vert
drdsdt  \notag \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right) -F\left(
x,y,z,r,s,t,f_{2}\left( v\right) \left( r,s,t\right) \right) \right\vert
drdsdt  \notag \\
&\leq l_{g}\left\vert h\left( u\right) \left( x,y,z\right) -h\left( v\right)
\left( x,y,z\right) \right\vert \notag \\
&+\int_{0}^{x}\int_{0}^{y}\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
l_{f_{1}}\left\vert f_{1}\left( u\right) \left( r,s,t\right) -f_{1}\left(
v\right) \left( r,s,t\right) \right\vert drdsdt \notag\\
&+\int_{0}^{\infty
}\int_{0}^{\infty }\int_{0}^{\infty }l_{F}\left( x,y,z,r,s,t\right)
l_{f_{2}}\left\vert f_{2}\left( u\right) \left( r,s,t\right) -f_{2}\left(
v\right) \left( r,s,t\right) \right\vert drdsdt \notag \\
&\leq l_{g}l_{h}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left( x+y+z\right)
}+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
l_{f_{1}}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left( r+s+t\right) }drdsdt
\notag \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left(
r+s+t\right) }drdsdt \\
&\leq l_{g}l_{h}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left( x+y+z\right)
}+l_{1}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left( x+y+z\right)
}+l_{2}\left\Vert u-v\right\Vert_{\tau } e^{\tau \left( x+y+z\right) }.
\end{align*}%
Then we have:%
\begin{equation*}
\left\Vert A\left( u\right) -A\left( v\right) \right\Vert_{\tau } \leq \left(
l_{g}l_{h}+l_{1}+l_{2}\right) \left\Vert u-v\right\Vert_{\tau }
\end{equation*}%
for all $u,v\in X_{\tau }.$ From $\left( C8\right) $ we have that $A$ is a
contraction. Hence $A$ is a $c-$Picard operator, with%
\begin{equation*}
c=\frac{1}{1-l_{g}l_{h}-l_{1}-l_{2}}.
\end{equation*}%
Hence the equation \eqref{eq1} has a unique solution in $X_{\tau }.$
\end{proof}

\section{Hyers-Ulam-Rassias stability}

In what follows we consider the equation%
\begin{align}\label{eq10}
&u\left( x,y,z\right) =g\left( x,y,z,h\left( u\right) \left( x,y,z\right)\right) \\
&+\int_{0}^{x}\!\!\! \int_{0}^{y}\!\!\! \int_{0}^{z}K\left( x,y,z,r,s,t,f_{1}\left(
u\right) \left( r,s,t\right) \right) drdsdt+ \int_{0}^{\infty
}\!\!\! \int_{0}^{\infty }\!\!\! \int_{0}^{\infty }F\left( x,y,z,r,s,t,f_{2}\left(
u\right) \left( r,s,t\right) \right) drdsdt \nonumber
\end{align}
and the inequality%
\begin{align}\label{eq20}
&\left\vert u\left( x,y,z\right) -g\left( x,y,z,h\left( u\right) \left(
x,y,z\right) \right) -\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right)
drdsdt-\right. \nonumber\\
&-\left.\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right)
drdsdt\right\vert \leq \varphi \left( x,y,z\right) ,
\end{align}
where $\left( E,\left\vert \cdot \right\vert \right) $ is a Banach space and
$\varphi \in C\left( \left[ 0,a\right) ^{3}, \mathbb{R}_{+}\right) $ is increasing$,g\in
C\left( \left[ 0,a\right) ^{3}\times E,E\right) ,K\in C\left( \left[
0,a\right) ^{6}\times E,E\right) ,F\in C\left( \left[ 0,a\right) ^{6}\times
E,E\right) ,h\in C\left( X_{\tau },X_{\tau }\right) ,f_{1}\in C\left(
X_{\tau },X_{\tau }\right) ,f_{2}\in C\left( X_{\tau },X_{\tau }\right) .$

\begin{theorem}
Under the conditions $\left( C1\right) -\left( C10\right) $  and
\begin{itemize}
\item[(i)]
there exists $N>0$ such that%
\begin{equation*}
\left\vert h\left( u\right) \left( x,y,z\right) -h\left( v\right) \left(
x,y,z\right) \right\vert \leq N\left\vert u\left( x,y,z\right) -v\left(
x,y,z\right) \right\vert ,\forall x,y,z\in \left[ 0,a\right) ,\forall u,v\in
X_{\tau };
\end{equation*}

\item[(ii)] $l_{g}N<1$,
\end{itemize}

if $u$ is a solution of \eqref{eq20} and $u^{\ast }$ is the unique
solution of \eqref{eq10}, we have%
\begin{equation*}
\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert
\leq C_{KFghf_{1}f_{2}}\varphi \left( x,y,z\right)
\end{equation*}%
where%
\begin{equation*}
C_{KFghf_{1}f_{2}}\varphi \left( x,y,z\right) =\frac{1}{1-l_{g}N}\exp \left(
\frac{m}{1-l_{g}N}\right) ,
\end{equation*}%
i.e the equation \eqref{eq10} is Hyers-Ulam-Rassias stable.
\end{theorem}

\begin{proof}
We have%
\begin{align*}
&\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert \\
&\leq \left\vert u\left( x,y,z\right) -g\left( x,y,z,h\left( u\right)
\left( x,y,z\right) \right) -\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right)
drdsdt\right.\\
&\left.-\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right)
drdsdt\right\vert  \\
&+\left\vert g\left( x,y,z,h\left( u\right) \left( x,y,z\right) \right)
-g\left( x,y,z,h\left( u^{\ast }\right) \left( x,y,z\right) \right)
\right\vert  \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}\left\vert K\left(
x,y,z,r,s,t,f_{1}\left( u\right) \left( r,s,t\right) \right) -K\left(
x,y,z,r,s,t,f_{1}\left( u^{\ast }\right) \left( r,s,t\right) \right)
\right\vert drdsdt \\
&+\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left\vert F\left(
x,y,z,r,s,t,f_{2}\left( u\right) \left( r,s,t\right) \right) -F\left(
x,y,z,r,s,t,f_{2}\left( u^{\ast }\right) \left( r,s,t\right) \right)
\right\vert drdsdt \\
&\leq \varphi \left( x,y,z\right) +l_{g}\left\vert h\left( u\right) \left(
x,y,z\right) -h\left( u^{\ast }\right) \left( x,y,z\right) \right\vert  \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
l_{f_{1}}\left\vert f_{1}\left( u\right) \left( r,s,t\right) -f_{1}\left(
u^{\ast }\right) \left( r,s,t\right) \right\vert drdsdt \\
&+\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\left\vert f_{2}\left( u\right) \left(
r,s,t\right) -f_{2}\left( u^{\ast }\right) \left( r,s,t\right) \right\vert
drdsdt.
\end{align*}%
From  conditions $\left( i\right) ,\left( ii\right) $ we have:%
\begin{align*}
&\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert
\leq \varphi \left( x,y,z\right) +l_{g}N\left\vert u\left( x,y,z\right)
-u^{\ast }\left( x,y,z\right) \right\vert  \\
&+\int_{0}^{x}\!\!\!\int_{0}^{y}\!\!\!\int_{0}^{z}l_{K}\left( x,y,z,r,s,t\right)
l_{f_{1}}\left\vert u\left( r,s,t\right) -u^{\ast }\left( r,s,t\right)
\right\vert drdsdt \\
&+\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\left\vert u\left( r,s,t\right) -u^{\ast }\left(
r,s,t\right) \right\vert drdsdt \\
&\leq \varphi \left( x,y,z\right) +l_{g}N\left\vert u\left( x,y,z\right)
-u^{\ast }\left( x,y,z\right) \right\vert \\
&+\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\!\!\!\int_{0}^{\infty}\left(l_{K}\left(
x,y,z,r,s,t\right) l_{f_{1}}+l_{F}\left( x,y,z,r,s,t\right) l_{f_{2}}\right)
\left\vert u\left( r,s,t\right) -u^{\ast }\left( r,s,t\right) \right\vert
drdsdt.
\end{align*}%
Then%
\begin{align*}
&\left( 1-l_{g}N\right) \left\vert u\left( x,y,z\right) -u^{\ast }\left(
x,y,z\right) \right\vert \\
&\leq \varphi \left( x,y,z\right) +\int_{0}^{\infty
}\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left( l_{K}\left(
x,y,z,r,s,t\right) l_{f_{1}}+l_{F}\left( x,y,z,r,s,t\right) l_{f_{2}}\right)
\left\vert u\left( r,s,t\right) -u^{\ast }\left( r,s,t\right) \right\vert
drdsdt
\end{align*}%
and we have%
\begin{align*}
&\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert
\leq \frac{\varphi \left( x,y,z\right) }{1-l_{g}N}\\
&+\frac{1}{1-l_{g}N}%
\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left(
l_{K}\left( x,y,z,r,s,t\right) l_{f_{1}}+l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\right) \left\vert u\left( r,s,t\right) -u^{\ast
}\left( r,s,t\right) \right\vert drdsdt.
\end{align*}%
From Wendorf Lemma \cite{FilatovSha} for unbounded domain it follows that%
\begin{align*}
&\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert\\
&\leq \frac{\varphi \left( x,y,z\right) }{1-l_{g}N}\exp \left[ \frac{1}{%
1-l_{g}N}\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\!\!\!\int_{0}^{\infty }\left(
l_{K}\left( x,y,z,r,s,t\right) l_{f_{1}}+l_{F}\left(
x,y,z,r,s,t\right) l_{f_{2}}\right) drdsdt\right]
\end{align*}%
and we have%
\begin{equation*}
\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert
\leq \frac{1}{1-l_{g}N}\exp \left[ \frac{m}{1-l_{g}N}\right] \cdot \varphi
\left( x,y,z\right)
\end{equation*}%
and%
\begin{equation*}
\left\vert u\left( x,y,z\right) -u^{\ast }\left( x,y,z\right) \right\vert
\leq C_{KFghf_{1}f_{2}}\cdot \varphi \left( x,y,z\right)
\end{equation*}%
where%
\begin{equation*}
C_{KFghf_{1}f_{2}}\varphi \left( x,y,z\right) =\frac{1}{1-l_{g}N}\exp \left(
\frac{m}{1-l_{g}N}\right) ,
\end{equation*}%
and the equation \eqref{eq10} is Hyers-Ulam-Rassias stable.
\end{proof}

\bibliographystyle{elsarticle-num}

\end{document}